OpenStudy (anonymous):

The monthly cost of driving a car depends on the number of miles driven. Lynn found that in May it cost her $418 to drive 320 mi and in June it cost her $594 to drive 760 mi. (a) Express the monthly cost C as a function of the distance driven, d, assuming that a linear relationship gives a suitable model.

OpenStudy (mertsj):

Two points: (320,418) and (760,595) Slope = 2/5 \[y-418=\frac{2}{5}(d-320)\]

OpenStudy (mertsj):

\[C(d)=\frac{2}{5}d-128+418\]

OpenStudy (mertsj):

\[C(d)=\frac{2}{5}d+290\]